A car travels at 40kph for 2 hrs, at 50 kph for 1 hr and at 20 kph for 0.5 hr, find the average speed of the car.


1
Expert's answer
2021-07-12T03:38:08-0400
"v_{av}=\frac{S}{T}\\v_{av}=\frac{(40)2+(50)1+(20)0.5}{2+1+0.5}=40\ kph"